OpenStudy (anonymous):

u=14i+10j & v=-15i-13j

OpenStudy (unklerhaukus):

two vectors, ...

OpenStudy (unklerhaukus):

Are you performing an operation with these vectors?

OpenStudy (anonymous):

subtract or add?

OpenStudy (anonymous):

i'm good with sub and add vectors

OpenStudy (anonymous):

if + then 14i+10j-15i-13j= -i-3j if - then 14i+10j+15i+13j=29i+23j